OpenStudy (anonymous):

Find the height of a cylinder with a radius of 5 cm and surface area of 1250 cm sq. A)13.42 B)26.83 C)34.79 D)20.0

OpenStudy (anonymous):

Use the formula :\[h = (SA - 2\pi r^2)/2\pi r\] You will have your answer...

OpenStudy (anonymous):

SO MY ANSWER WOULD BE (C) ?

OpenStudy (anonymous):

Yes your answer will be C.. you are damn true...